The legs of a triangle are 5 and 12 inches. What is the hypotenuse of the triangle?
Answer by jim_thompson5910(35256)

a^2 + b^2 = c^2

5^2 + 12^2 = c^2

25 + 144 = c^2

169 = c^2

c^2 = 169

c = sqrt(169)

c = 13

The hypotenuse is 13 inches.